OpenStudy (anonymous):

Catherina conducted a survey of four school cafeterias to find the number of students who like burgers for lunch. The results of her survey are recorded in the table below: School Cafeteria SurveySchool Total Number of Students in the Cafeteria Number of Students Who Like Burgers A 35 15 B 42 13 C 21 8 D 41 16 Which school has the greatest percentage of students who like burgers for lunch? School A School B School C School D

OpenStudy (anonymous):

@phi @sleepyhead314 @Elsa213 @micahm Are any of you available to help me please?

OpenStudy (micahm):

the first one openstudy.com/updates/55671d33e4b05d462cd5361d

OpenStudy (anonymous):

You positive its A?

OpenStudy (micahm):

i just gather the site I am not an answering person for math i not that good but this site should help you here is the sec one openstudy.com/updates/55671d33e4b05d462cd5361d openstudy.com/updates/558ab39be4b09ad72e9661ac

OpenStudy (anonymous):

Alright Thank you (: i have one more can i tag you for it?

OpenStudy (dustinrathke24):

15/35=.429 13/42=.310 8/21=.381 16/41=.390 School A has the highest percentage of students who like burgers for lunch.